OpenStudy (stamp):

Use implicit differentiation to find a formula for dy dx. \[xy^2-3ye^x=x-3\]

OpenStudy (stamp):

\[xy^2-3ye^x=x-3\]

OpenStudy (stamp):

been a while since I've done implicit differentiation so I'm a bit hazy

OpenStudy (loser66):

just derivative both sides

OpenStudy (stamp):

\[d(xy^2-3ye^x)=1\]

OpenStudy (loser66):

(y(xy-3e^x))'=1, right? do as usual

OpenStudy (loser66):

hey, I lost hihihihi... please take steps

OpenStudy (stamp):

brb I need water/air haha

OpenStudy (loser66):

\[y'(xy-3e^x)+y(xy-3e^x)'=1\] \[(y'xy -y'3e^x)+y(y+xy'-3e^x)=1\] \[y'xy -y'3e^x+y^2+yxy'-y3e^x=1\] factor y' \[y'(2xy-3e^x)+y^2-y3e^x =1\] \[y'=\dfrac{1-y^2+3ye^x}{2xy-3e^x}\]

OpenStudy (stamp):

Alright I'm gonna try and replicate that without looking on my paper, then look to see if I did the right thing

OpenStudy (loser66):

good guy,

OpenStudy (stamp):

Your answer is correct, and I'm close to it all clicking. I'm going to close this question because the rest is on me and I can feel it, but thank you.

OpenStudy (stamp):

\[x d(y^2)+y^2d(x)-3yd(e^x)-e^xd(3y)=d(x-3)\]\[y'2xy+y^2-3ye^x-y'3e^x=1\]\[y'(2xy-3e^x)+y^2-3ye^x=1\]\[y'(2xy-3e^x)=1-y^2+3ye^x\]\[y'=\frac{1-y^2+3ye^x}{2xy-3e^x}\]